# GITA: Graph to Visual and Textual Integration for Vision-Language Graph Reasoning Project Page: v-graph.github.io

### v4

@@ 5.4 Case Study
In this section, we present representative examples of graph information provided in both visual and textual formats, which may offer some intuitive interpretations for our experimental results. Specifically, we highlight a case where the GITA-7B (VO) model outperforms its LLMs-based counterpart in Figure 2 (a), and its opposite scenario in Figure 2 (b).
The task depicted in Figure 2 (a) is cycle detection and the correct answer is ‘No’. This is successfully identified by the vision-only GITA-7B model, while the text-based Vicuna-7B fails to recognize it. In this instance, cycle patterns appear to be more readily recognizable in visual graphs, whereas text-based LLMs face the challenge of identifying cycles from disordered textual descriptions of edges, which could inherently involve greater complexity and more challenges.
@@
On the other hand, the fixed layout of visual graphs presented in GVLQA-BASE may impede the visual encoder in identifying the shortest path between two nodes, although we have verified layout augmentation can greatly improve the graph reasoning abilities of models, as shown in Sec. 5.2. This limitation might arise from the confusion caused by the visual distance within an image, without considering the weights between the nodes. For instance, in Figure 2(b), the correct answer is ’4->6->0’, which visually appears as a more convoluted path but numerically has a shorter path length of $3=1+2$. In contrast, the incorrect answer given by GITA-7B (vision-only) is ’4->2->0’, which has a higher path length cost of $4=1+3$ but visually seems like a more direct shortcut. This observation further validates the effectiveness of employing layout augmentation to enhance performance in this task. Layout variations of visual graphs play a crucial role in mitigating the visual confusion caused by the spatial arrangement within a visual graph. However, it seems more effective for text-based LLMs to handle explicitly separated nodes and weights, as illustrated in the red-highlighted text in Figure 2(b).
@@ 6 Conclusion
In this paper, we propose an end-to-end framework, called GITA, for vision-language-based graph reasoning. Our extensive experiments validate the superiority of incorporating visual information into instruction-based graph reasoning. Furthermore, we conduct a comparative analysis of the four proposed visual graph augmentations and identify layout augmentation as the most effective approach for enhancing visual graphs. This finding offers valuable insights for the development of larger-scale datasets aimed at facilitating vision-language-based graph reasoning. Lastly, we highlight the potential application value of the proposed GVLQA dataset as a pretrained dataset.
Appendix A Characteristics of GITA
Generalizability. By effectively separating task specifications from graph structures and employing distinct components to manage them individually, GITA can be utilized to address a diverse set of graph reasoning tasks without necessitating alterations to its underlying architecture. In this way, when facing different types of tasks, the only adjustment required is to modify the instructions in the questioner without modifying other components.
Flexibility. The components of GITA offer a broad spectrum of choices: (i) The graph visualizer seamlessly integrates various graphic tools such as graphviz, matplotlib and networkx. (ii) The questioner can be either manual-template crafted or automatically generated by an agent. (iii) The VLM reasoner is compatible with almost all the existing VLMs. Such versatility empowers the implementation of GITA to be highly flexible.
User-friendliness. By utilizing existing VLM models as the reasoner, GITA inherits their language capabilities, allowing it to respond in natural languages.
Scalability on Large Graphs.
In real-world scenarios, the graphs can be very large. However, feeding the entire graph into the language model is not feasible due to the limited context length of text tokens. Additionally, visually representing such large graphs can lead to congestion and difficult to interpret. To address those challenges and maintain the usability for both text and visual inputs, GITA samples subgraphs around the target nodes or links within $k$ hops (Sec 3.2). This has been commonly used in graph-based learning Kipf & Welling (2016); Hamilton et al. (2017); Xu et al. (2018); Jui et al. (2023); Velickovic et al. (2017); Chen et al. (2022), and ensures efficiency without compromising usability.
Appendix B Datasets Statistics
@@
Appendix C Illustrations for Visualization Tools in Graph Visualizer
@@
Appendix D Graph-describing Templates
@@
Appendix E Examples of Manual-template-based and LLM-agent-bootstrapped Query Generation
@@
LLM-agent-bootstrapped Query Generation. Figure 5 presents an example of employing a bootstrapped LLM agent, such as ChatGPT(OpenAI, 2022), for monster-hunting gaming. By incorporating task-specific information into the prompt, including node/edge meanings and task responsibilities, the LLM agent automatically generates a response that serves as the desired task-specific query. Compared to using manual templates, bootstrapping LLM agents for automated synthesis is more flexible and economic as it can take advantage of LLM agents to automatically interpret context and generate appropriate queries for various scenarios and minimize manual effort with changing conditions. Such properties make it suitable for dynamic or bespoke tasks, such as robotic planning or complex gaming scenarios.
@@
Appendix F Experiment Settings
@@
Detailed Settings for GVLQA Dataset During the evaluation, the temperature is set to 0 for all baselines. All fine-tuning experiments are conducted on an NVIDIA DGX station with 8×A100 GPUs. We split the GVLQA dataset in the ratio of 7:3 for training and testing, respectively. The accuracy (%) metrics are computed by comparing the prediction and ground truths with exact matching. We use the next-token-prediction loss to fine-tune the LoRA (Hu et al., 2021) adapters of LLMs and the vision-to-text projector. Visual graphs are encoded as visual embeddings by a visual encoder. Visual embeddings are concatenated with the embeddings of textual descriptions and instructions (i.e., questions), then fed to the text decoder to generate the answer.
@@
Appendix G Illustrations of GVLQA subsets

### v5

@@ 5.5 Case Study
In this section, we present examples of graph information provided in both visual and textual formats, which offer some intuitive interpretations for our experimental results. Figure 2 (a) shows an example where the GITA-7B (VO) model outperforms its LLMs-based counterpart, and Figure 2 (b) shows an opposite scenario.
The task depicted in Figure 2(a) is cycle detection and the correct answer is ‘No’. This is predicted successfully by the vision-only GITA-7B model, while the text-based Vicuna-7B fails. In this example, recognizing cycle patterns is much easier in visual graphs, whereas text-based LLMs struggle with disordered textual descriptions of edges, which could inherently involve greater complexity and more challenges.
On the other hand, the fixed layout of visual graphs presented in GVLQA-BASE may impede the visual encoder in identifying the shortest path between two nodes, although we have verified layout augmentation can greatly improve the graph reasoning abilities of models, as shown in Sec. 5.2. This limitation might arise from the confusion caused by the visual distance within an image, without considering the weights between the nodes. For instance, in Figure 2(b), the correct answer is ’4->6->0’, which visually appears as a more convoluted path but numerically has a shorter path length of $3=1+2$. In contrast, the incorrect answer given by GITA-7B (vision-only) is ’4->2->0’, which has a higher path length cost of $4=1+3$ but visually seems like a more direct shortcut. This observation further validates the effectiveness of employing layout augmentation to enhance performance in this task. Layout variations of visual graphs play a crucial role in mitigating the visual confusion caused by the spatial arrangement within a visual graph. However, it seems more effective for text-based LLMs to handle explicitly separated nodes and weights, as illustrated by the text (in red) in Figure 2(b).
@@ 6 Conclusion
In this paper, we propose an end-to-end framework called GITA for vision-language-based graph reasoning. Extensive experiments validate the superiority of incorporating visual information into instruction-based graph reasoning. Furthermore, we conduct comparative analysis of the four proposed visual graph augmentations and identify layout augmentation as the most effective approach for enhancing visual graphs. This finding offers valuable insights for the development of larger-scale datasets aimed at facilitating vision-language-based graph reasoning. Lastly, we highlight the potential application value of the proposed GVLQA dataset as a pretrained dataset.
Acknowledgements
This work was supported by NSFC key grant 62136005 and NSFC general grant 62076118, and in part by the Research Grants Council of the Hong Kong Special Administrative Region (Grants 16200021 and 16202523).
Appendix A Visual Modality Enhances Effectiveness by Uncovering Critical Substructures
In this section, we present a case study to highlight the complementary role of the visual modality in graph reasoning tasks. The visual modality excels at recognizing beneficial substructures or local patterns, which are crucial for effective graph reasoning. For instance, identifying the "hop number" is essential for shortest path calculations, recognizing "leaf nodes" is vital for topological sorting, and detecting "cycles" is necessary to avoid in Hamilton path construction. We extracted these substructures from the GVLQA-Base dataset and manually labeled them. By employing a frozen Vision Transformer (ViT) in the LLaVA framework with a trainable Multi-Layer Perceptron (MLP) decoder, we achieved identification accuracies of 89.92%, 95.16%, and 92.39% for hop number counting, leaf node identification, and cycle detection, respectively. In contrast, using a pre-trained BERT model with the same trainable MLP decoder resulted in significantly lower accuracies of 55.47%, 26.33%, and 60.32% for the same tasks. Therefore, the enhanced effectiveness of integrating visual and textual modalities can be attributed to the additional structural information provided by the visual modality, which facilitates the identification of these critical substructures.
Appendix B Advantages of GITA Over Traditional Graph Neural Networks
GITA offers several advantages over traditional Graph Neural Networks (GNNs) in terms of generalizability, flexibility, and user-friendliness:
Unlike GNNs, which require task-specific feature engineering and architecture adjustments, GITA employs a unified model architecture for all tasks, demonstrating its generalizability. By separating task specifications from graph structures, GITA can handle various graph reasoning tasks seamlessly. Additionally, it exhibits strong zero-shot capabilities, allowing it to perform well on tasks it has not been explicitly trained on, which is a feature not commonly found in traditional GNNs.
Besides, traditional GNNs often demand specialized knowledge in model architectures and coding to accommodate diverse tasks, posing a challenge for non-experts. In contrast, GITA overcomes this barrier by employing language-based templates for task adaptation, enhancing its flexibility. This flexibility enables GITA to effectively handle a broad spectrum of tasks, offering a framework that can be customized to specific requirements using daily language, without the necessity of profound expertise in graph neural networks.
Moreover, by leveraging existing VLMs, GITA can respond in natural language, allowing for intuitive graph reasoning with simple queries like "Is there a cycle in this graph?" This stands in contrast to the unreadable vector representations typically used in GNNs, significantly enhancing GITA’s user-friendliness.
Appendix C Datasets Statistics
@@
Table 7: Average numbers of nodes and edges for each task in GVLQA.
Average / Task Connect Cycle TS SP MaxFlow BGM HP
#nodes 25.01 23.42 21.86 13.65 13.90 21.13 13.24
#edges 95.46 23.66 114.10 23.99 49.16 51.03 45.05
@@
Appendix D Illustrations for Visualization Tools in Graph Visualizer
@@
Appendix E Graph-describing Templates
@@
Appendix F Examples of Manual-template-based and LLM-agent-bootstrapped Query Generation
@@
LLM-agent-bootstrapped Query Generation. Figure 5 presents an example of employing a bootstrapped LLM agent, such as ChatGPT[48], for monster-hunting gaming. By incorporating task-specific information into the prompt, including node/edge meanings and task responsibilities, the LLM agent automatically generates a response that serves as the desired task-specific query. Compared to using manual templates, bootstrapping LLM agents for automated synthesis is more flexible and economic as it can take advantage of LLM agents to automatically interpret context and generate appropriate queries for various scenarios and minimize manual effort with changing conditions. Such properties make it suitable for dynamic or bespoke tasks, such as robotic planning or complex gaming scenarios.
@@
Appendix G Experiment Settings
@@
Detailed Settings for GVLQA Dataset During the evaluation, the temperature is set to 0 for all baselines. All fine-tuning experiments are conducted on an NVIDIA DGX station with 8×A100 GPUs. We split the GVLQA dataset in the ratio of 7:3 for training and testing, respectively. The accuracy (%) metrics are computed by comparing the prediction and ground truths with exact matching. We use the next-token-prediction loss to fine-tune the LoRA [28] adapters of LLMs and the vision-to-text projector. Visual graphs are encoded as visual embeddings by a visual encoder. Visual embeddings are concatenated with the embeddings of textual descriptions and instructions (i.e., questions), then fed to the text decoder to generate the answer.
@@
Appendix H Illustrations of GVLQA subsets
@@ Appendix I Limitation
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [NA] .
  • [NA] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[NA] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS paper checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Please refer to the abstract part and the contribution enumeration at the tail of the introduction part.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Please refer to Appendix I
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [NA]
  14. Justification: The paper does not involve any theoretical result.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We show fundamental experiment settings in Section 5, and more details for experiments settings in Appendix G. Besides, we provide the complete codes as supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The complete codes are included, and the proposed GVLQA dataset is released with common access.
  25. Guidelines:
    • The answer NA means that the paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experiment details are given in both Section 5 and Appendix G.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: The paper does not include error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the machine (type and storage) requirements in Appendix G.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We make sure the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [NA]
  49. Justification: The research does not have concerns about societal impacts because it is designed for general-purpose graph reasoning.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: The paper includes using an graph visualizer to generate abstract graph images, however, these images are focus on graph structure, without any sensitive information.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited necessary assets and conduct CC-BY 4.0 for our codes and datasets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The code and other supplementary materials are followed with readme and instructions.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [NA]
  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [NA]
  74. Justification: the paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.